OpenStudy (anonymous):

Five trigonometric functions of theta. Cos0=3/5,Tan0<0. Please explain.

OpenStudy (jdoe0001):

\(\bf cos(\theta)=\cfrac{3}{5}\qquad tan(\theta)<0 \\ \quad \\ tan(\theta)<0\Leftarrow\textit{another way of saying, tangent is negative} \\ \quad \\ {\color{blue}{ tan(\theta)=\cfrac{sin(\theta)}{cos(\theta)}\qquad -tan(\theta)=\cfrac{sin(\theta)}{-cos(\theta)}\qquad -tan(\theta)=\cfrac{-sin(\theta)}{cos(\theta)}}} \\ \quad \\ cos(\theta)=\cfrac{3}{5}\implies \cfrac{adjacent}{hypotenuse}\implies \cfrac{a=3}{c=5}\qquad \textit{what's "b"?} \\ \quad \\ c^2=a^2+b^2\implies \pm \sqrt{c^2-a^2}={\color{blue}{ b}}\qquad recall\implies sin(\theta)=\cfrac{{\color{blue}{ b}}}{c}\) so... the pythagorean theorem doesn't tell us if the root will be +/- so what is "b"? well, notice above, we know the tangent is negative for that to happen either the cosine or sine has to be negative we know the cosine isn't, so the sine will be, thus the "b" component is negative

OpenStudy (jdoe0001):

so once you find "b", you can pretty much use all three, "a", "b" and "c" to get the trig functions values

OpenStudy (anonymous):

Would b=4???

OpenStudy (jdoe0001):

recall, the tangent is negative, the cosine is positive, or "a" is positive, so "b" have to be negative \(\bf c^2=a^2+b^2\implies \sqrt{c^2-a^2}={\color{blue}{ b}}\implies \pm\sqrt{25-9}={\color{blue}{ b}}\iff -4\)

OpenStudy (anonymous):

Thanks but where would five trigonometric functions come from??

OpenStudy (jdoe0001):

http://htmartin.myweb.uga.edu/6190/resources/Trig%20Functions.gif <--- this one looks better

OpenStudy (anonymous):

Thank you!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Just for clarification sine would be -0.8

OpenStudy (jdoe0001):

well, you'd usually keep it as fraction, but I gather you could use the decimal format as well, \(\bf sin(\theta)=\cfrac{b=-4}{c=5}\implies -\cfrac{4}{5}\iff -0.8\)

OpenStudy (anonymous):

How you would you find the exact six trigonometric function at point (-2,6)???